# Risk-Averse Learning with Delayed feedback

### v1

@@ 3.2 Two-point risk-averse learning
Selecting $\delta=T^{-\frac{a}{4}}$ and $\eta=(T+D)^{-\frac{1}{2}}$, Algorithm 1 achieves ${\rm R}_{2}(T)=\tilde{\mathcal{O}}\Big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{1}{2}}\Big{)}$ with probability $1-2\gamma$. $\blacksquare$
Remark 3.
Theorems 1 and 2 show that Algorithm 2 based on two-point gradient estimate achieves better performance than Algorithm 1 based on one-point gradient estimate, for all $a>0$. The explanation follows. Let $\delta_{i}$ and $\eta_{i}$, for $i=1,2$, represent the smoothing parameter and the step size of Algorithm 1 and 2, respectively. When $a\geq 1$, we have ${\rm R}_{1}(T)=\tilde{\mathcal{O}}\big{(}(T+D)^{\frac{3}{4}}\big{)}$, ${\rm R}_{2}(T)=\tilde{\mathcal{O}}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{1}{2}}\big{)}$. The regret bound of Algorithm 1 is worse than Algorithm 2 since $(T+D)^{\frac{3}{4}}\geq T^{1-\frac{a}{4}}+(T+D)^{\frac{1}{2}}$. When $0<a<\frac{2}{3}$, for the third and fifth terms of (3.1), we have $\frac{\eta_{1}L_{0}dU}{\delta_{1}}D+\frac{\eta_{1}}{2}\frac{d^{2}U^{2}}{\delta_{1}^{2}}T=\tilde{\mathcal{O}}\big{(}(T+D)^{1-\frac{3a}{4}}T^{\frac{a}{4}}\big{)}$, as in Remark 1. Similarly, $\frac{\eta_{2}L_{0}dU}{\delta_{2}}D+\frac{\eta_{2}}{2}\frac{d^{2}U^{2}}{\delta_{2}^{2}}T=\tilde{\mathcal{O}}\big{(}(T+D)^{\frac{1}{2}}\big{)}$. The comparison statement also holds since $(T+D)^{1-\frac{3a}{4}}T^{\frac{a}{4}}\geq(T+D)^{\frac{1}{2}}$. When $\frac{2}{3}<a<1$, for the forth term of (3.1), we have $\frac{D_{x}^{2}}{\eta_{1}}=\tilde{\mathcal{O}}\big{(}(T+D)^{\frac{3a}{4}}\big{)}$. Similarly, we have $\frac{D_{x}^{2}}{\eta_{2}}=\tilde{\mathcal{O}}\big{(}(T+D)^{\frac{1}{2}}\big{)}$ and $(T+D)^{\frac{3a}{4}}\geq(T+D)^{\frac{1}{2}}$. Thus, we conclude that the upper bound of ${\rm R}_{2}(T)$ is tighter than that of ${\rm R}_{1}(T)$.
Remark 4.
Let $D\leq c_{1}T^{\alpha}$, for $c_{1}$, $\alpha>0$. By following the proof of Proposition 1 and Corollary 1, we have ${\rm R}_{2}(T)=\tilde{\mathcal{O}}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{\alpha}{2}}\big{)}$. When $d_{t}=T-t$, meaning none feedback arrives over the horizon $T$, we have $\tilde{\mathcal{O}}\big{(}D\big{)}=T^{2}$. It implies that $\alpha$ is bounded by $\alpha\leq 2$ naturally, and the two-point algorithm has a high tolerance for delay.
@@
In this section, we consider the dynamic pricing problem for parking lots, see [22]. Factors such as parking prices, availability, and locations generally influence driving decisions. This encourages us to dynamically adjust the parking price according to real-time demand. Denote $r_{t}\in[0,1]$ as curb occupancy rate. The occupancy rate is influenced by the price $x_{t}$ and environmental uncertainties $\xi_{t}$:
$r_{t}=\xi_{t}+Ax_{t},$
where $A=-0.15$ is the estimated price elasticity, which is determined by [22] through analyzing the real-world data. Assume that uncertainty $\xi$ obeys a uniform distribution between $[0.9,1.1]$, i.e., $\mathcal{U}$. To make it easy to find a parking space, it is desirable to maintain an occupancy rate of $70\%$. Moreover, environmental changes such as weather and holidays affect the desired occupancy rate. We choose the desired occupancy as $r_{d}=0.7+\sin{\frac{\pi}{1000}t}$. The loss function is defined as
@@
where $\nu=0.001$ is the regularization parameter. To prevent either overcrowding or insufficient occupancy, our goal is to minimize the risk-averse objective function
[…]
where the risk level is selected as $\alpha=0.5$. We set the initial price value as $x_{0}=1$ and restrict the potential prices to $[1,5]$.
We choose the maximum delay as $\bar{d}=200$ and let the delay obey a uniform distribution with the range of $[0,200]$. Set the sample number as $n_{t}=8$ for all $t$. Shaded areas represent $\pm$ one standard deviation over $10$ runs. Fig. 1 depicts parking prices and the corresponding occupancies generated by Algorithms 1 and 2. Fig. 1 shows that the prices and occupancies generated by Algorithm 2 better track the desired values compared to the Algorithm 1. Fig. 2 depicts CVaR values and cumulated loss $\sum_{t=1}^{T}C_{t}(\hat{x}_{t})$ and $\sum_{t=1}^{T}\big{(}C_{t}(\hat{x}_{t}^{1})+C_{t}(\hat{x}_{t}^{2})\big{)}/2$ generated by Algorithms 1 and 2, respectively. Fig. 2 shows that Algorithm 2 achieves lower cumulated CVaR values than Algorithm 1, i.e., the two-point risk-averse learning algorithm outperforms the one-point algorithm. Moreover, both Figs. 1 and 2 reveal that Algorithm 2 yields results with a lower variance compared to Algorithm 1. This suggests that Algorithm 2 produces more consistent and stable outcomes than Algorithm 1.
Figure 1: From top to bottom: parking prices generated by the one-point algorithm (Algorithm 1) and by the two-point algorithm (Algorithm 2); the corresponding resulting occupancies and the desired occupancy rate $r_{d}$.
@@
Fig. 3 depicts the resulting occupancies and the regret achieved by two-point risk-averse learning algorithm with maximum delay $\bar{d}=\{0,200,400\}$. The sampling number is set as $n_{t}=8$. Fig. 3 shows that, as the delay increases, the variance of the resulting occupancy and the cumulated loss both increase. Fig. 4 depicts the resulting occupancies and the regret achieved by Algorithm 2 with sampling number $n_{t}=\{1,2,16\}$ for all $t$, while fixing the maximum delay as $\bar{d}=200$. Fig. 4 shows that, as the sampling number increases, the variance of the resulting occupancy and the cumulated loss both decrease.
Figure 3: The occupancies and cumulated losses generated by Algorithm 2 under maximum individual delays $\bar{d}=\{0,200,400\}$.
@@
5 CONCLUSIONS
This paper investigated risk-averse learning with delayed feedback, where the delays are varying but bounded. We reorganized the arriving feedback into virtual time slots based on their reception order. Both one-point and two-point zeroth-order optimization approaches are utilized to estimate the CVaR gradient, with function values being queried multiple times under perturbed actions in both cases. We then developed corresponding one-point and two-point risk-averse learning algorithms based on these CVaR gradient estimates. The regret bounds of these algorithms are analyzed in terms of cumulative delay and total sampling number. The theoretical results indicate that the two-point risk-averse learning algorithm outperforms the one-point algorithm. Furthermore, the one-point algorithm achieves sublinear regret when the cumulative and individual delays meet specific bound conditions, while the two-point algorithm requires only minimal restrictions on both cumulative and individual delays.
@@ 6 Appendix
Proof. We start from the first claim. Denote $L_{\mu(\tau)-1}=m$, which means the learner receives $m$ batches feedback before $t_{1}=\mu(\tau)$. Due to the presence of delay, we have $10\leq m\leq\mu(\tau)-$ 1. The learner receives the $\tau$-th batch feedback at slot $t_{2}=\mu(\tau)+d_{\mu(\tau)}\geq t_{1}$. In other words, the learner receives at least $m$ batches feedback before $t_{2}$. Then, we have $\tau\geq m+1$, and further that $\tilde{s}_{\tau}\geq m+1-1-m=0$.
@@
with $\mathcal{S}=\{\tau|1\leq\tau\leq L_{T},\mu(\tau)>L_{T}-\bar{d}\}$. The last equality of (6) follows from separating $\sum_{\tau=1}^{L_{T}}L_{\mu(\tau)-1}$ into two parts: $L_{\mu(\tau)-1}$ for $\tau\in\{s|1\leq s\leq L_{T},\mu(s)\in\{1,\dots,L_{T}-\bar{d}\}\}$, i.e., those generated during $t\in\{1,\dots,L_{T}-\bar{d}\}$, and the remaining, i.e., $L_{\mu(\tau)-1}$, for $\tau\in\mathcal{S}$. For the first and the second term of (6), we have
@@
where the first equality is from $L_{0}=0$. We next explain the inequality. In the absence of delay, the learner receives $t$ batches of feedback by the end of slot $t$. Then, $\sum_{t=1}^{{L_{T}-\bar{d}-1}}t$ can be interpreted as each packet is accounted once at each slot since its arrival over the horizon $L_{T}-\bar{d}-1$. Accordingly, $\sum_{t=1}^{{L_{T}-\bar{d}-1}}L_{t}$ can be interpreted as each packet is accounted once at each slot since its arrival in the presence of delay. By Assumption 1, the losses $J_{t}(\cdot)$, for $t\in\{1,\dots,L_{T}-2\bar{d}-1\}$, arrive by end of $L_{T}-\bar{d}-1$ slot. Moreover, $J_{t}(\cdot)$ is undercounted $d_{t}$ times by slot $t+\bar{d}$. Thus, we have $\sum_{t=1}^{L_{T}-2\bar{d}-1}t-\sum_{t=1}^{L_{T}-\bar{d}-1}L_{t}\leq\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}$. For the last term of (6), we have

### v2

@@ 3.2 Two-point risk-averse learning
Selecting $\delta=T^{-\frac{a}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{1}{2}}\sqrt{\bar{d}P_{T}}$, Algorithm 2 achieves ${\rm DR}_{2}(T)=\tilde{\mathcal{O}}\bigg{(}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{1}{2}}\big{)}\sqrt{\bar{d}P_{T}}\bigg{)}$ with probability $1-2\gamma$. $\blacksquare$
Remark 4
By comparing the regret bounds of Algorithms 1 and 2, i.e., $\mathrm{DR}_{1}(T)$ and $\mathrm{DR}_{2}(T)$ in Theorem 1 and 2, it can be observed that the two-point risk-averse learning algorithm (Algorithm 2) outperforms the one-point risk-averse learning algorithm (Algorithm 1), by achieving a smaller regret bound, for all $a>0$. Additionally, by comparing the order of the cumulative delay $D$, one finds that Algorithm 2 has a higher tolerance for delay compared to Algorithm 1.
@@ 4 Simulation
In this section, we consider the dynamic pricing problem for parking lots, see [28]. Factors such as parking prices, availability, and locations generally influence driving decisions. This encourages us to dynamically adjust the parking price according to real-time demand. Denote $r_{t}\in[0,1]$ as curb occupancy rate. The occupancy rate is influenced by the price $x_{t}$ and environmental uncertainties $\xi_{t}$: $r_{t}=\xi_{t}+Ax_{t}$, where A=0.15A=-0.15 is the estimated price elasticity, which is determined by [28] through analyzing the real-world data. Assume that the uncertainty $\xi$ follows a uniform distribution between $[0.9,1.1]$, i.e., $\mathcal{U}$. Generally, it is easier to find a parking space when the occupancy rate is maintained at $70\%$. Thus, we modeled the desired occupancy as $r_{d}=0.7$. To prevent either overcrowding or insufficient occupancy, our goal is to minimize the risk-averse objective function
[…]
where the risk level is selected as $\alpha=0.5$ and the loss function is given by
@@
where $\nu=0.001$ is the regularization parameter. We set the initial price value as $x_{0}=1$ and restrict the potential prices to the range $[1,5]$.
Figure 3: From top to bottom: parking prices generated by the one-point algorithm (Algorithm 1) and by the two-point algorithm (Algorithm 2); the corresponding resulting occupancies and the desired occupancy rate $r_{d}=0.7$.
Figure 4: From top to bottom: the CVaR values under the price $\hat{x}_{t}$ generated by Algorithm 1 and under the prices $\hat{x}_{t}^{1}$ and $\hat{x}_{t}^{2}$ generated by Algorithm 2; the cumulated CVaR value achieved by the one-point algorithm (Algorithm 1) and the two-point algorithm (Algorithm 2).
Figure 5: From top to bottom: parking prices generated by the one-point algorithm (Algorithm 1) and by the two-point algorithm (Algorithm 2); the corresponding resulting occupancies and the desired occupancy rate $r_{d}=0.6$ when $t\in[0,200]$ and $r_{d}=0.7$ when $t\in(200,500]$.
@@
We assume that the parking lot periodically monitors the occupancy rate and adjusts prices accordingly. Assume that the unit of time for an iteration is one week, which means the parking lot observes the occupancy rate and changes prices every week. However, due to data aggregation or computational overhead, these observations may be delayed. Assume that the random delays obey a uniform distribution with the range of $[0,5]$, with a maximum delay of $\bar{d}=5$. Namely, the maximum information delay is $5$ weeks. Set the sample number as $n_{t}=8$ for all $t$. Shaded areas represent $\pm$ one standard deviation over $20$ runs. Fig. 3 depicts parking prices and the corresponding occupancies generated by Algorithms 1 and 2. Fig. 3 demonstrates that the prices and occupancies generated by Algorithm 2 track the desired values faster compared to Algorithm 1. Fig. 4 depicts CVaR values and the dynamic regret $\mathrm{DR}(T)$ generated by Algorithms 1 and 2, respectively. Fig. 4 shows that the two-point risk-averse learning algorithm outperforms the one-point algorithm by achieving lower regret.
Additionally, we consider the scenario where the desired occupancy changes over time due to variations in demand or usage patterns. Assume that the desired occupancy rate is $r_{d}=0.65$ when $t\in[0,200]$ and $r_{d}=0.7$ when $t\in(200,500]$. Fig. 5 depicts parking prices and the corresponding occupancies generated by Algorithms 1 and 2, which demonstrates that the prices and occupancies generated by Algorithm 2 track desired values faster compared to Algorithm 1. Fig. 6 depicts CVaR values and the dynamic regret $\mathrm{DR}(T)$ of Algorithms 1 and 2 when facing the changing desired occupancy, respectively. Fig. 6 shows that the two-point risk-averse learning algorithm outperforms the one-point algorithm by achieving a lower regret. Moreover, Figs. 3-6 reveal that Algorithm 2 yields results with a lower variance compared to Algorithm 1. This suggests that Algorithm 2 produces more consistent and stable outcomes than Algorithm 1.
Figure 7: The occupancies and cumulated losses generated by Algorithm 2 under maximal individual delays $\bar{d}=\{0,20,50\}$.
@@
Fig. 7 depicts the resulting occupancies and the regret achieved by the two-point risk-averse learning algorithm under the maximum delay of $\bar{d}=\{0,20,50\}$. The sampling number is set as $n_{t}=8$. Fig. 7 shows that, the occupancy under Algorithm 2 converges to the desired value slower as the delay increases, and the variance of the resulting occupancy and the cumulated loss both increase. This effect is particularly pronounced when $\bar{d}=50$.
In Fig. 8, we fix the maximal delay as $\bar{d}=5$ and depict the resulting occupancies and the regret of Algorithm 2 with sampling number $n_{t}=\{1,2,16\}$ for all $t$. Fig. 8 shows that, as the sampling number increases, the occupancy under Algorithm 2 converges to the desired value faster and the variance of the resulting occupancy and the cumulated loss both decrease.
5 Conclusion
This paper investigated risk-averse learning with delayed feedback, where the delays are random yet bounded. We developed the one-point and two-point risk-averse learning algorithms, where the one-point and two-point zeroth-order optimization approaches are employed to estimate the CVaR gradient, respectively. In both cases, function values are queried multiple times under perturbed actions to improve the accuracy of the CVaR gradient estimate. To capture the error induced by using outdated feedback to process the gradient descent, we sorted the arriving feedback according to their arrival order and calculated the cumulative delay. The dynamic regret bounds of these algorithms are then analyzed in terms of the cumulative delay, the total sampling number and the path length of comparators. The theoretical results demonstrated that both the one-point and two-point risk-averse learning algorithms achieve sublinear regrets. Furthermore, the two-point risk-averse learning algorithm outperforms the one-point algorithm by achieving a smaller regret bound.
@@ 6 Appendix
Proof of Lemma 4. The first claim is from [17], we provide its proof here for completeness. Denote $L_{\mu(\tau)-1}=m$, which means that the learner receives $m$ batches feedback before $t_{1}=\mu(\tau)$. Due to the delay, we have $0\leq m\leq\mu(\tau)-1$. The learner receives the $\tau$-th batch feedback at slot $t_{2}=\mu(\tau)+d_{\mu(\tau)}\geq t_{1}$. In other words, the learner receives at least $m$ batches feedback before $t_{2}$. Then, we have $\tau\geq m+1$, and further that $\tilde{s}_{\tau}\geq m+1-1-m=0$. Additionally, before the slot $\mu(\tau)$, the learner receives $L_{\mu(\tau)-1}$ batches feeback. Note that $L_{\mu(\tau)-1}\geq\mu(\tau)-\bar{d}-1$, we have $\tilde{s}_{\tau}=\tau-1-L_{\mu(\tau)-1}\leq\tau-1-t(\tau)+1+\bar{d}\leq 2\bar{d}.$
@@
with $\mathcal{S}=\{\tau|1\leq\tau\leq L_{T},\mu(\tau)>L_{T}-\bar{d}\}$. For the first and the second terms of (6), we have
@@
where the first equality is from $L_{0}=0$. We next explain the inequality. In the absence of delay, the learner receives $t$ batches of feedback by the end of slot $t$. Then, $\sum_{t=1}^{{L_{T}-\bar{d}-1}}t$ can be interpreted as each packet is accounted once at each slot since its arrival over the horizon $L_{T}-\bar{d}-1$. Accordingly, $\sum_{t=1}^{{L_{T}-\bar{d}-1}}L_{t}$ means that each packet is accounted once at each slot since its arrival in the presence of delay. Additionally, due to the delay, each $J_{t}(\cdot)$ is undercounted $d_{t}$ times by slot $t+\bar{d}$. By Assumption 1, the losses $J_{t}(\cdot)$, for $t\in\{1,\dots,L_{T}-2\bar{d}-1\}$, arrive before the end of $L_{T}-\bar{d}-1$ slot. Thus, we have $\sum_{t=1}^{L_{T}-2\bar{d}-1}t-\sum_{t=1}^{L_{T}-\bar{d}-1}L_{t}\leq\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}$. For the last term of (6), we have
@@
Delays result in out-of-order feedback, which imposes challenges in comparing the decisions generated by the algorithm against the optimal ones. The following lemma is prepared for the regret analysis in Theorems 1 and 2.
Lemma 5
Consider the virtual mapping $\tilde{x}_{\tau}=x_{\mu(\tau)}$, the following inequalities hold.
$\sum_{\tau=2}^{L_{T}}\|\tilde{x}_{\tau-1}^{\ast}-\tilde{x}_{\tau}^{\ast}\|$ $\leq 2\bar{d}P_{T}$ (49)
$\sum_{\tau=2}^{L_{T}}\|x_{\mu(\tau)}^{\ast}-\tilde{x}_{\tau}^{\ast}\|$ $\leq 2\bar{d}P_{T}$ (50)
Proof of Lemma 5. For some $\mu(\tau)=t$, we have $t-\bar{d}-1\leq\mu(\tau-1)\leq t+\bar{d}-1$. Then, for $t=\mu(\tau)>\bar{d}+1$, we have
$\hskip 10.00002pt\|\tilde{x}_{\tau-1}-\tilde{x}_{\tau}\|=\|x_{\mu(\tau-1)}-x_{\mu(\tau)}\|$
$\leq\max_{i\in\{t-\bar{d}-1,\dots,t+\bar{d}-1\}}\|x_{i}-x_{t}\|\leq\sum_{k=t-\bar{d}}^{t+\bar{d}-1}\|x_{k-1}-x_{k}\|.$
For $1\leq\mu(\tau)<\bar{d}+1$, we have $1\leq\mu(\tau-1)\leq t+\bar{d}-1$. Then, we have
$\|\tilde{x}_{\tau-1}-\tilde{x}_{\tau}\|\hskip-1.99997pt\leq\hskip-1.99997pt\max_{i\in\{1,\dots,t+\bar{d}-1\}}\|x_{i}-x_{t}\|\hskip-1.99997pt\leq\hskip-1.99997pt\sum_{k=2}^{t+\bar{d}-1}\|x_{k-1}-x_{k}\|,$
Furthermore, $\mu(\tau)\leq T$. Thus, we obtain
$\sum_{\tau=2}^{L_{T}}\|\tilde{x}_{\tau-1}^{\ast}-\tilde{x}_{\tau}^{\ast}\|$ $\leq\sum_{t=2}^{T}\sum_{k=\min\{2,t-\bar{d}\}}^{t+\bar{d}-1}\|x_{k-1}^{\ast}-x_{k}^{\ast}\|\leq 2\bar{d}P_{T}$ (51)
The next is to prove (50). To analyze the range of $\mu(\tau)$, we consider the following two extreme cases. First, consider a real-time slot $t=k$. Under the assumption of a maximum delay of $\bar{d}$, at most $\bar{d}-1$ pieces of feedback have not been received. More specifically, suppose that the feedback $C_{k}$ is received by the learner without delay, while the feedbacks $\{C_{k-\bar{d}+1},\dots,C_{k-1}\}$ are all delayed by $\bar{d}$ time slots. In this case, for $\tau=k-\bar{d}+1$, we have $\mu(\tau)=k$. This yields $\mu(\tau)\geq\tau-\bar{d}+1$. Second, consider the opposite scenario where $C_{t}$ is delayed by $\bar{d}$ step, and thus received at time $t+\bar{d}$. By then, the learner has received at most $k+\bar{d}-1$ feedbacks. In this case, for $\tau=k+\bar{d}$, we have $\mu(\tau)=k$. This yields $\mu(\tau)\leq\tau+\bar{d}$. Combining both bounds, we obtain that $\mu(\tau)\in[\tau-\bar{d}+1,\tau+\bar{d}]$ and $x_{\mu(\tau)}\in\{x_{\tau-\bar{d}+1},\dots,x_{\tau+\bar{d}}\}$. Following similar arguments, one can prove $\tilde{x}_{\tau}\in\{x_{\tau-\bar{d}+1},\dots,x_{\tau+\bar{d}}\}$. Therefore, we have
$\hskip 10.00002pt\sum_{\tau=2}^{L_{T}}\|x_{\mu(\tau)}^{\ast}-\tilde{x}_{\tau}^{\ast}\|\leq\sum_{\tau=2}^{L_{T}}\max_{i,j\in\{\tau-\bar{d}+1,\tau+\bar{d}\}}\|x_{i}^{\ast}-x_{j}^{\ast}\|$
$\leq\sum_{\tau=2}^{L_{T}}\sum_{k=\min\{2,\tau-\bar{d}\}}^{\tau+\bar{d}}\|x_{k-1}^{\ast}-x_{k}^{\ast}\|\leq 2\bar{d}P_{T}$ (52)
where the last equality is from Definition 1. $\blacksquare$